OpenStudy (anonymous):

help me to the derivative (cos2x + sin2x) / (cos2x - sin2x)

OpenStudy (anonymous):

derivative of tan (pi/4+2x) 2sec^2 (pi/4+2x)

OpenStudy (anonymous):

I do not understand your solution

OpenStudy (anonymous):

\[\frac {\cos2x +\sin2x}{\cos2x-\sin2x}= \frac {1+\tan2x}{1-\tan2x}\]

OpenStudy (anonymous):

now do u understand??

OpenStudy (anonymous):

I want to resolve it with the derivative and differentiation rules

OpenStudy (anonymous):

hello @maribor76 didn't u understand my approach??

OpenStudy (anonymous):

the second post of me have u understood??

OpenStudy (anonymous):

\[\frac{\text{Cos}[2 x]+\text{Sin}[2 x]}{\text{Cos}[2 x]-\text{Sin}[2 x]}\text{=}\text{ Csc}\left[\frac{\pi }{4}-2 x\right] \text{Sin}\left[\frac{\pi }{4}+2 x\right] \]The derivative of the RHS is\[2 \text{Cos}\left[\frac{\pi }{4}+2 x\right] \text{Csc}\left[\frac{\pi }{4}-2 x\right]+2 \text{Cot}\left[\frac{\pi }{4}-2 x\right] \text{Csc}\left[\frac{\pi }{4}-2 x\right] \text{Sin}\left[\frac{\pi }{4}+2 x\right] \]The above simplifed is:\[-\frac{4}{\sin (4 x)-1} \]

OpenStudy (anonymous):

\[\frac {1+\tan2x}{1-\tan2x}= \tan(\pi/4 +2x)\]\[\frac d{dx} \tan(\frac {\pi}4+2x) = 2\sec^2(\frac {\pi}4+2x)\]

OpenStudy (anonymous):

Good work dipankarstudy!

OpenStudy (anonymous):

Thank you all now I get it